# Dispersive regime of multiphoton qubit-oscillator interactions

### v1

@@ Abstract
The dispersive regime of $n$-photon qubit-oscillator interactions is analyzed using Schrieffer-Wolff perturbation theory. Effective Hamiltonians are derived up to the second order in the perturbation parameters. These effective descriptions reveal higher-order qubit-oscillator cross-Kerr and oscillator self-Kerr terms. The cross-Kerr term combines a qubit Pauli operator with an $n$-degree polynomial in the oscillator photon number operator, while the self-Kerr term is an $(n-1)$-degree polynomial in the oscillator photon number operator. In addition to the higher-order Kerr terms, a qubit-conditional $2n$-photon squeezing term appears in the effective non-rotating-wave-approximation Hamiltonian. Furthermore, perturbation theory is applied to the case of multiple qubits coupled to a shared oscillator. A photon-number-dependent qubit-qubit interaction emerges in this case, which can be leveraged to tune the effective multiqubit system parameters using the oscillator state. Results for the converse setup of multiple oscillators and a single qubit are also derived. In this case, a qubit-conditional oscillator-oscillator nonlinear interaction is found. The effective descriptions developed here offer a simple and intuitive physical picture of dispersive multiphoton qubit-oscillator interactions that can aid in the design of implementations harnessing their desired nonlinear effects.
@@ I Introduction
The aforementioned studies spurred an active area of research focused on the spectral and dynamical features of the generalized Rabi and JC models. One particularly peculiar phenomenon occurs in some variants of the $n$-photon Rabi models; the feature known as the “spectral collapse” where the joint qubit-oscillator spectra and eigenstates transition from a discretely indexed set to a degenerate continuum above a critical qubit-oscillator coupling strength [14, 15]. For the two-photon Rabi model, a quasi-analytical solution was found in the studies of Refs. [16, 17] based on the same method used by Braak [18] to solve the one-photon case. Other works used generalized Bogoliubov transformations to arrive at the same solution [19, 20]. In parallel with the spectral and analytical studies of multiphoton models, there have been a number of quantum information applications developed that leverage nonlinear qubit-oscillator interactions [21, 22, 23]. Experimentally feasible implementations were proposed for superconducting circuits [22, 24, 25, 26] and trapped ions [21, 27], and very recently, qubit-conditional $n$-photon squeezing for $n=2,3,4$ was experimentally realized in a trapped ions setup [23].
@@
In this paper, we aim to add more insight to the study of multiphoton qubit-oscillator interactions. While quasi-analytical solutions for the two-photon Rabi model were found [16, 17, 19, 20], they lack an intuitive and physically-motivated picture. We seek to uncover a more intuitive picture than the rigorous mathematical solutions of the spectra of $n$-photon qubit-oscillator models in the dispersive regime, where the qubit-oscillator $n$-photon coupling is much smaller than the $n$-photon qubit-oscillator detuning as we will define later. We resort to second-order Schrieffer-Wolff (SW) perturbation theory to obtain analytic expressions for the qubit-oscillator spectra in the dispersive regime by generalizing techniques employed with the Rabi and JC models [28, 29, 30]. We then consider two further multipartite generalizations of this multiphoton dispersive perturbation theory; the case of many qubits nonlinearly coupled to a shared oscillator and the converse case of many oscillators nonlinearly coupled to a shared qubit.
@@
The paper is structured as follows. In Sec. II, we review relevant previous work. In particular, we recall the typically used SW perturbation theory for the dispersive regime of linear qubit-oscillator interactions that we seek to generalize. Section III defines the multiphoton dispersive regime and lays out the generalized SW transformations for the RWA and non-RWA regimes. The analytically-obtained spectra are compared to numerical results to establish a regime of validity. The transformations are then used in the multiqubit-single-oscillator scenario in Sec. IV. The multioscillator-single-qubit converse case is discussed in Sec. V. The results are summarized and conclusions are given in Sec. VI. Appendix A outlines the details of the combinatorics and commutators used in the SW perturbation theory developed in the main text.
@@ II Background
In this section, we briefly introduce the Rabi model and its subsequent approximations and relevant generalizations that are instrumental for the results to be presented. We begin with models relying on linear interactions in the case of a single qubit-oscillator system followed by the multipartite cases of many qubits linearly coupled to a single oscillator and many oscillators linearly coupled to a single qubit. In each case, we derive an effective RWA and non-RWA dispersive Hamiltonian using second-order perturbation theory. Finally, we introduce the multiphoton generalized models of interest for which we seek to develop a similar perturbation theory for the RWA and non-RWA $n$-photon dispersive regimes.
@@ II.1 Single qubit-oscillator system
where $\chi=g^{2}/\Delta$ is the dispersive shift. Here, we employ second-order SW perturbation theory since the first-order term in the BCH expansion, $\lambda[\hat{H}_{\text{JC}},\hat{X}_{-}]$, yields terms quadratic in $\lambda$, and the second-order term, $\lambda^{2}[[\hat{H}_{\text{JC}},\hat{X}_{-}],\hat{X}_{-}]/2!$, leads to terms that are quadratic and quartic in $\lambda$. Thus, we need to go to second order in SW perturbation theory to account for all the terms that are quadratic in $\lambda$. This will be the case for all perturbative expansions in this paper. This Hamiltonian is typically employed for quantum-nondemolition measurements of a qubit using an oscillator. The perturbative second-order expansion is typically considered valid when $\lambda\ll 1$. However, strictly speaking, there is an additional constraint; the oscillator average occupation must remain lower than the critical photon number $n_{c}=\lambda^{2}/4=g^{2}/4\Delta^{2}$. Thus, the approximate description becomes less accurate for larger $\lambda$ and/or higher oscillator photon occupation. In either of these cases, higher-order corrections in the expansion are required [28]. The critical photon number places a bound on the linear dispersive regime where the oscillator does not exhibit any nonlinearity, e.g. $(\hat{a}^{\dagger}\hat{a})^{2}$, as it would from higher-order corrections. There have been works studying the higher-order corrections and the effect of the qubit-induced oscillator nonlinearity on the efficacy of qubit readout [32].
@@ II.2 Multiple qubits coupled to a single oscillator
$\,\,\,\,\,\,\,\,+\sum_{k>l}\widetilde{\chi}^{(l,m)}(\hat{\sigma}_{+}^{(l)}\hat{\sigma}_{-}^{(m)}+\hat{\sigma}_{-}^{(l)}\hat{\sigma}_{+}^{(m)})$ (12)
@@
$\,\,\,\,\,\,\,\,+\sum_{k>l}(\widetilde{\chi}^{(l,m)}-\widetilde{\xi}^{(l,m)})(\hat{\sigma}_{x}^{(l)}\hat{\sigma}_{x}^{(m)}),$ (13)
where $\xi^{(l)}=(g^{(l)})^{2}/\Sigma^{(l)}$, and $\widetilde{\xi}^{(l,m)}=g^{(l)}g^{(m)}\left(\frac{1}{\Sigma^{(l)}}+\frac{1}{\Sigma^{(m)}}\right).$ The presence of the counter-rotating terms gives rise to an Ising-type mediated qubit-qubit interaction, $\hat{\sigma}_{x}^{(l)}\hat{\sigma}_{x}^{(m)}$, which is not excitation preserving.
@@ II.3 Multiple oscillators coupled to a single qubit
In this regime, the qubit-conditional interactions include both a beamsplitter and two-mode squeezing interaction (non-excitation preserving). The presence of both terms allows for different continous variable gates such as a qubit-controlled SUM gate or a qubit-controlled-control-$Z$ gate [49, 43].
@@ III Single Qubit-Oscillator Multiphoton Interaction Dispersive Regime
$\hat{H}_{n-\text{R}}=\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}\hat{\sigma}_{x}(\hat{a}^{\dagger n}+\hat{a}^{n}),$ (18)
@@
$\hat{H}_{n-\text{R}}\simeq\hat{H}_{n-\text{JC}}=\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}(\hat{\sigma}_{+}\hat{a}^{n}+\hat{\sigma}_{-}\hat{a}^{\dagger n}).$ (20)
@@
The $n$R model for $n>2$ suffers from spectral instabilities due to the unboundedness of the Hamiltonian from below which leads to states with infinite negative energies at any nonzero coupling ($g_{n}>0$). This means that calculations even for low coupling strengths will be plagued by the infinite negative energies affecting the Hamiltonian spectrum. Other studies have shown that the $n$R model for $n>2$ is, strictly speaking, not self adjoint and, thus, they concluded it is unphysical [51]. It is important to note that the $n$JC Hamiltonian while exactly solvable also suffers from negative infinite energies. This becomes obvious when expressing the $n$JC Hamiltonian in the following form:
@@
where the $2\times 2$ matrices act on the subspaces spanned by {|$\{\ket{e,l},\ket{g,l+n}\}$ for nonnegative integers $l.$ This form explicitly shows the aforementioned singlets and doublets which are represented by the first and second terms, respectively. To show the negative infinite energies, we can simply obtain the doublet energies by diagonalize the $2\times 2$ matrix in each fixed $l$ subspace;
@@
For very large $l$, the second term (under the square root) scales as $l^{n/2}$ while the first term which is strictly postive, scales as $l$. This means that there are infinitely many states with infinite negative energies for all $n>2$. Even though the $n$JC Hamiltonian is exactly solvable, it is also problematic because it is unbounded from below. However, when studying the $n$R or $n$JC Hamiltonian, one must keep in mind that, in practice, these models never arise as the sole terms in a realistic system Hamiltonian. Rather, they are arrived at as an effective description in a weak-coupling and low-energy regime. These effective descriptions typically neglect other terms that bound these spectral instabilities such as unbounded states (see, for example, the derivations of effective multiphoton Hamiltonians in Refs [22] and [52], where spurious terms that make the total system Hamiltonian bounded are neglected to obtain similar effective descriptions).
@@
$\hat{\widetilde{H}}_{n-\text{R}}=$ $\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}\hat{\sigma}_{x}(\hat{a}^{\dagger n}+\hat{a}^{n})$
@@ III.1 RWA dispersive regime
$\hat{H}_{\text{Disp,RWA}}^{(n)}$ $=\hat{U}_{\text{Disp,RWA}}^{(n)\dagger}\hat{H}_{n-\text{JC}}\hat{U}_{\text{Disp,RWA}}^{(n)}$
$=\hat{H}_{n-\text{JC}}+\lambda_{n}[\hat{H}_{n-\text{JC}},\hat{X}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n-\text{JC}},\hat{X}_{-n}],\hat{X}_{-n}]+...$
@@
The Hamiltonian contains two polynomials in the photon number operator, $\chi_{n}\sum_{k=0}^{n}C_{n,k}^{+}\hat{\sigma}_{z}(\hat{a}^{\dagger}\hat{a})^{k}$ and $\chi_{n}\sum_{k=0}^{n-1}C_{n,k}^{-}(\hat{a}^{\dagger}\hat{a})^{k}$, with the qubit-state-dependent polynomial truncating at order $n$ and the other truncating at order $n-1$. The qubit-state-independent polynomial truncating at order $n-1$ has to do with the coefficients $C_{n,k}^{-}$; they are zero when $n=k$ for all $n$. Due to this fact, the qubit-state-dependent series causes a larger shift between energy levels. The qubit-state dependence is reflected in the sign of the shift; the shift is negative when the qubit is in |$\ket{g}$, and it is positive when the qubit is in |$\ket{e}$. Interestingly, for interaction orders greater than $n=2$, the qubit-oscillator interaction induces a dispersive anharmonicity in the oscillator. This usually occurs for the dispersive linear interaction ($n=1$) when accounting for expansion terms beyond second order, and due to their higher-order nature, the strength of such terms is usually very small [54]. For a given order $n$, the qubit-state-dependent $n$th order shift, as well as the lower-order dispersive shifts ($n-1$, $n-2$, …), are all present in the Hamiltonian. Interestingly, the lower-order shifts are always larger than the $n$th order shift as seen in the table of $C_{n,k}^{+}$ values; $C_{n,n}^{+}=2$ for all $n$. The resulting higher-order self-Kerr terms induce multiphoton blockades in the presence of a drive on the oscillator [55]. Additionally, when the oscillator is initialized in a coherent state, these higher-order self-Kerr terms can be used to make multi-component cat states [56].
Figure 1: Energy-level dispersive shifts for the $j$th Fock state. The qubit-state-independent shifts in the RWA (green) and non-RWA (orange) cases are always weaker than the qubit-state-dependent RWA (blue) and non-RWA (red) shifts. This due to the fact that the former’s polynomial is of degree $n-1$, while the polynomial of the latter is of degree $n$. Additionally, both RWA shifts are larger than the non-RWA shifts since $\chi_{n}>\xi_{n}$.

### v2

@@ Abstract
The dispersive regime of $n$-photon qubit-oscillator interactions is analyzed using Schrieffer-Wolff perturbation theory. Effective Hamiltonians are derived up to the second order in the perturbation parameters. These effective descriptions reveal higher-order qubit-oscillator cross-Kerr and oscillator self-Kerr terms. The cross-Kerr term combines a qubit Pauli operator with an $n$-degree polynomial in the oscillator photon number operator, while the self-Kerr term is an $(n-1)$-degree polynomial in the oscillator photon number operator. In addition to the higher-order Kerr terms, a qubit-conditional $2n$-photon squeezing term appears in the effective non-rotating-wave-approximation Hamiltonian. Furthermore, perturbation theory is applied to the case of multiple qubits coupled to a shared oscillator. A photon-number-dependent qubit-qubit interaction emerges in this case, which can be leveraged to tune the effective multiqubit system parameters using the oscillator state. Results for the converse setup of multiple oscillators and a single qubit are also derived. In this case, a qubit-conditional oscillator-oscillator nonlinear interaction is found. The spectral instabilities plaguing multiphoton qubit-oscillator models are carefully treated by introducing stabilizing higher-order terms in the Hamiltonian. The stabilizing terms preserve low-photon subspaces, avoid negative infinite energies and facilitate reliable numerical calculations used to validate analytical predictions. The effective descriptions developed here offer a simple and intuitive physical picture of dispersive multiphoton qubit-oscillator interactions that can aid in the design of implementations harnessing their desired nonlinear effects.
@@ I Introduction
The aforementioned studies spurred an active area of research focused on the spectral and dynamical features of the generalized Rabi and JC models. One particularly peculiar phenomenon occurs in some variants of the $n$-photon Rabi models; the feature known as the “spectral collapse” where the joint qubit-oscillator spectra and eigenstates transition from a discretely indexed set to a degenerate continuum above a critical qubit-oscillator coupling strength [14, 15]. For the two-photon Rabi model, a quasi-analytical solution was found in the studies of Refs. [16, 17] based on the same method used by Braak [18] to solve the one-photon case. Other works used generalized Bogoliubov transformations to arrive at the same solution [19, 20]. In parallel with the spectral and analytical studies of multiphoton models, there have been a number of quantum information applications developed that leverage nonlinear qubit-oscillator interactions [21, 22, 23]. Experimentally feasible implementations were proposed for superconducting circuits [22, 24, 25, 26] and trapped ions [21, 27]. Generally, the engineering of nonlinear interactions between a qubit and an oscillator remains an active area of research with novel nonlinear interactions realized in various platforms such as opto/electromechanics [28, 29] and trapped ions [23].
@@
In this paper, we aim to add more insight to the study of multiphoton qubit-oscillator interactions. While quasi-analytical solutions for the two-photon Rabi model were found [16, 17, 19, 20], they lack an intuitive and physically-motivated picture. We seek to uncover a more intuitive picture than the rigorous mathematical solutions of the spectra of $n$-photon qubit-oscillator models in the dispersive regime, where the qubit-oscillator $n$-photon coupling is much smaller than the $n$-photon qubit-oscillator detuning as we will define later. We resort to second-order Schrieffer-Wolff (SW) perturbation theory to obtain analytic expressions for the qubit-oscillator spectra in the dispersive regime by generalizing techniques employed with the Rabi and JC models [30, 31, 32]. We then consider two further multipartite generalizations of this multiphoton dispersive perturbation theory; the case of many qubits nonlinearly coupled to a shared oscillator and the converse case of many oscillators nonlinearly coupled to a shared qubit. Furthermore, we discuss the spectral instabilities arising in multiphoton qubit-oscillator interaction models and remedy them using stabilizing higher-order terms in the Hamiltonian.
@@
The paper is structured as follows. In Sec. II, we review relevant previous work. In particular, we recall the typically used SW perturbation theory for the dispersive regime of linear qubit-oscillator interactions that we seek to generalize. The reader who is familiar with SW perturbation theory applied to the linear interaction can skip Sec. II. Section III defines the multiphoton dispersive regime and lays out the generalized SW transformations for the RWA and non-RWA regimes. The analytically-obtained spectra are compared to numerical results to establish a regime of validity. The transformations are then used in the multiqubit-single-oscillator scenario in Sec. IV. The multioscillator-single-qubit converse case is discussed in Sec. V. The results are summarized and conclusions are given in Sec. VI. Appendix A outlines the details of the combinatorics and commutators used in the SW perturbation theory developed in the main text. Appendix B explains the spectrum stabilization relying on higher-order terms and its details.
@@ II Background
In this section, we review the Rabi model and its subsequent approximations and relevant generalizations that are instrumental for the results to be presented. We begin with models relying on linear interactions in the case of a single qubit-oscillator system followed by the multipartite cases of many qubits linearly coupled to a single oscillator and many oscillators linearly coupled to a single qubit. In each case, we derive an effective RWA and non-RWA dispersive Hamiltonian using second-order perturbation theory.
@@ II.1 Single qubit-oscillator system
where $\chi=g^{2}/\Delta$ is the dispersive shift. Here, we employ second-order SW perturbation theory since the first-order term in the BCH expansion, $\lambda[\hat{H}_{\text{JC}},\hat{X}_{-}]$, yields terms quadratic in $\lambda$, and the second-order term, $\lambda^{2}[[\hat{H}_{\text{JC}},\hat{X}_{-}],\hat{X}_{-}]/2!$, leads to terms that are quadratic and quartic in $\lambda$. Thus, we need to go to second order in SW perturbation theory to account for all the terms that are quadratic in $\lambda$. This will be the case for all perturbative expansions in this paper. This Hamiltonian is typically employed for quantum-nondemolition measurements of a qubit using an oscillator. The perturbative second-order expansion is typically considered valid when $\lambda\ll 1$. However, strictly speaking, there is an additional constraint; the oscillator average occupation must remain lower than the critical photon number $n_{\text{ph},c}=1/(4\lambda^{2})=\Delta^{2}/4g^{2}$ [30, 34]. The critical photon number stems from the JC energy eigenvalues,
$E_{\pm,n_{\text{ph}}}^{\text{JC}}=\left(n_{\text{ph}}+\frac{1}{2}\right)\omega_{o}\pm\sqrt{g^{2}(n_{\text{ph}}+1)+\Delta^{2}},$ (7)
with $n_{\text{ph}}$ being the photon occupation. The dispersive description makes the approximation
$\sqrt{g^{2}(n_{\text{ph}}+1)+\Delta^{2}}$ $=\Delta\sqrt{\lambda^{2}(n_{\text{ph}}+1)+1}$
$\simeq\Delta\left(\frac{\lambda^{2}}{2}(n_{\text{ph}}+1)+1\right)$
$=\frac{\chi}{2}(n_{\text{ph}}+1)+\Delta.$ (8)
Thus, when the energy difference between adjacent dressed states becomes comparable to the coupling strength, i.e. $g\sqrt{n_{\text{ph}}}\sim|\Delta|$, the linearization of the square root term, $\sqrt{g^{2}(n_{\text{ph}}+1)+\Delta^{2}}$, is no longer a valid approximation. Equivalently, this occurs when the photon occupation $n_{\text{ph}}\sim|\Delta|^{2}/g^{2}.$ Thus, the approximate description becomes less accurate for larger $\lambda$ and/or higher oscillator photon occupation. In either of these cases, higher-order corrections in the expansion are required [30]. The critical photon number places a bound on the linear dispersive regime where the oscillator does not exhibit any nonlinearity, e.g. $(\hat{a}^{\dagger}\hat{a})^{2}$, as it would from higher-order corrections. Higher-order corrections and the effect of qubit-induced oscillator nonlinearity have been studied in work on the efficacy of qubit readout [34].
@@ II.2 Multiple qubits coupled to a single oscillator
$\,\,\,\,\,\,\,\,+\sum_{m>l}\widetilde{\chi}^{(l,m)}(\hat{\sigma}_{+}^{(l)}\hat{\sigma}_{-}^{(m)}+\hat{\sigma}_{-}^{(l)}\hat{\sigma}_{+}^{(m)})$ (14)
@@
$\,\,\,\,\,\,\,\,+\sum_{m>l}(\widetilde{\chi}^{(l,m)}-\widetilde{\xi}^{(l,m)})(\hat{\sigma}_{x}^{(l)}\hat{\sigma}_{x}^{(m)}),$ (15)
where $\xi^{(l)}=(g^{(l)})^{2}/\Sigma^{(l)}$, and $\widetilde{\xi}^{(l,m)}=g^{(l)}g^{(m)}\left(\frac{1}{\Sigma^{(l)}}+\frac{1}{\Sigma^{(m)}}\right).$ The presence of the counter-rotating terms gives rise to an Ising-type mediated qubit-qubit interaction, $\hat{\sigma}_{x}^{(l)}\hat{\sigma}_{x}^{(m)}$, which is not excitation preserving. A related study, in which the coupling strength was pushed to the deep-strong coupling regime, was performed in Ref. [44].
@@ II.3 Multiple oscillators coupled to a single qubit
In this regime, the qubit-conditional interactions include both a beamsplitter and two-mode squeezing interaction (non-excitation preserving). The presence of both terms allows for different continous variable gates such as a qubit-controlled SUM gate or a qubit-controlled-control-$Z$ gate [52, 46]. The MMR model was leveraged in the dispersive regime to generate entangled cat states in two three-dimensional superconducting cavities mediated by a transmon qubit [53].
@@ III Single Qubit-Oscillator Multiphoton Interaction Dispersive Regime
$\hat{H}_{n\text{R}}=\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}\hat{\sigma}_{x}(\hat{a}^{\dagger n}+\hat{a}^{n}),$ (20)
@@
$\hat{H}_{n\text{R}}\simeq\hat{H}_{n\text{JC}}=\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}(\hat{\sigma}_{+}\hat{a}^{n}+\hat{\sigma}_{-}\hat{a}^{\dagger n}).$ (22)
@@
The $n$R model for $n>2$ suffers from spectral instabilities due to the unboundedness of the Hamiltonian from below which leads to states with infinite negative energies at any nonzero coupling ($g_{n}>0$). This means that calculations even for low coupling strengths will be plagued by the infinite negative energies affecting the Hamiltonian spectrum. Other studies have shown that the $n$R model for $n>2$ is, strictly speaking, not self adjoint and, thus, they concluded it is unphysical [55]. It is important to note that the $n$JC Hamiltonian, while exactly solvable, also suffers from negative infinite energies. This becomes obvious when expressing the $n$JC Hamiltonian in the following form:
@@
where the $2\times 2$ matrices act on the subspaces spanned by {|$\{\ket{e,l},\ket{g,l+n}\}$ for nonnegative integers $l.$ This form explicitly shows the aforementioned singlets and doublets which are represented by the first and second terms, respectively. To show the negative infinite energies, we can simply obtain the doublet energies by diagonalizing the $2\times 2$ matrix in each fixed $l$ subspace;
@@
For very large $l$, the second term (under the square root) scales as $l^{n/2}$ while the first term which is strictly postive, scales as $l$. This means that there are infinitely many states with infinite negative energies for all $n>2$. Even though the $n$JC Hamiltonian is exactly solvable, it is also problematic because it is unbounded from below. However, when studying the $n$R or $n$JC Hamiltonian, one must keep in mind that, in practice, these models never arise as the sole terms in a realistic system Hamiltonian. Rather, they are arrived at as an effective description in a weak-coupling and low-energy regime. These effective descriptions typically neglect other terms that bound these spectral instabilities such as unbounded states (see, for example, the derivations of effective multiphoton Hamiltonians in Refs. [22] and [56], where spurious terms that make the total system Hamiltonian bounded are neglected to obtain similar effective descriptions).
@@
$\hat{\widetilde{H}}_{n\text{R}}=$ $\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}\hat{\sigma}_{x}(\hat{a}^{\dagger n}+\hat{a}^{n})$
@@ III.1 RWA dispersive regime
$\hat{H}_{\text{Disp,RWA}}^{(n)}$ $=\hat{U}_{\text{Disp,RWA}}^{(n)\dagger}\hat{H}_{n\text{JC}}\hat{U}_{\text{Disp,RWA}}^{(n)}$
$=\hat{H}_{n\text{JC}}+\lambda_{n}[\hat{H}_{n\text{JC}},\hat{X}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n\text{JC}},\hat{X}_{-n}],\hat{X}_{-n}]+...$
@@
The Hamiltonian contains two polynomials in the photon number operator, $\chi_{n}\sum_{k=0}^{n}C_{n,k}^{+}\hat{\sigma}_{z}(\hat{a}^{\dagger}\hat{a})^{k}$ and $\chi_{n}\sum_{k=0}^{n-1}C_{n,k}^{-}(\hat{a}^{\dagger}\hat{a})^{k}$, with the qubit-state-dependent polynomial truncating at order $n$ and the other truncating at order $n-1$. The qubit-state-independent polynomial truncating at order $n-1$ has to do with the coefficients $C_{n,k}^{-}$; they are zero when $n=k$ for all $n$. Due to this fact, the qubit-state-dependent series causes a larger shift between energy levels. The qubit-state dependence is reflected in the sign of the shift; the shift is negative when the qubit is in |$\ket{g}$, and it is positive when the qubit is in |$\ket{e}$. Interestingly, for interaction orders greater than $n=2$, the qubit-oscillator interaction induces a dispersive anharmonicity in the oscillator. This usually occurs for the dispersive linear interaction ($n=1$) when accounting for expansion terms beyond second order, and due to their higher-order nature, the strength of such terms is usually very small [58]. For a given order $n$, the qubit-state-dependent $n$th order shift, as well as the lower-order dispersive shifts ($n-1$, $n-2$, …), are all present in the Hamiltonian. Interestingly, the lower-order shifts are always larger than the $n$th order shift as seen in the table of $C_{n,k}^{+}$ values; $C_{n,n}^{+}=2$ for all $n$.
The resulting higher-order self-Kerr terms induce multiphoton blockades in the presence of a drive on the oscillator [59]. Additionally, when the oscillator is initialized in a coherent state, these higher-order self-Kerr terms can be used to make multi-component cat states [60]. It is interesting to note that the one-photon JC Hamiltonian under exact diagonalization reveals higher-order Kerr terms with similar combinatorial factors [61].
Figure 1: Energy-level dispersive shifts for the $j$th Fock state. The bare frequency of the $j$th Fock is $j\omega_{o}$, and in the presence of a dispersive multiphoton interaction with the qubit, second-order SW perturbation theory predicts that the Fock state experiences a shift with four contributions; two qubit-state-dependent shifts and two qubit-state-independent shifts. The shifts are also classified according the interaction terms giving rise to them; RWA, $\hat{\sigma}_{+}\hat{a}^{n}$ and $\hat{\sigma}_{-}\hat{a}^{\dagger n}$, and non-RWA, $\hat{\sigma}_{+}\hat{a}^{\dagger n}$ and $\hat{\sigma}_{-}\hat{a}^{n}$, terms. The qubit-state-independent shifts in the RWA (green) and non-RWA (orange) cases are always weaker than the qubit-state-dependent RWA (blue) and non-RWA (red) shifts. This is due to the fact that the former’s polynomial is of degree $n-1$, while the polynomial of the latter is of degree $n$. Additionally, both RWA shifts are larger than the non-RWA shifts since $\chi_{n}>\xi_{n}$.
Figure 2: Critical photon number scaling for different interaction orders. The plots show the function $(\Delta_{n}/g_{n})^{2/n}$ for different $\Delta_{n}$ values and orders $n.$ Energies below the curve are considered well-approximated by the RWA dispersive description. Overall, the permissible region for states with larger energy content shrinks for higher order $n.$ As expected, increasing $\Delta_{n}$ improves the approximation (as $\lambda_{n}$ becomes smaller), and, thus, increases the area under the curve.